Show that the transformation x = z −"\frac{b}{3a}" converts the cubic equation ax3 + bx2 + cx + d= 0 into one of the form z3 + 3Hz + G = 0, i.e. the x2 term goes away.


1
Expert's answer
2021-11-30T16:23:47-0500
"x=z-\dfrac{b}{3a}"

"x^2=z^2-\dfrac{2bz}{3a}+\dfrac{b^2}{9a^2}"

"x^3=z^3-\dfrac{bz^2}{a}+\dfrac{b^2z}{3a^2}-\dfrac{b^3}{27a^3}"

Substitute


"az^3-bz^2+\dfrac{b^2z}{3a}-\dfrac{b^3}{27a^2}"

"+bz^2-\dfrac{2b^2z}{3a}+\dfrac{b^3}{9a^2}"

"cz-\dfrac{bc}{3a}+d=0"

"az^3+(-\dfrac{b^2}{3a}+c)z+(\dfrac{2b^3}{9a^2}-\dfrac{bc}{3a}+d)=0"

"z^3+3(-\dfrac{b^2}{3a^2}+\dfrac{c}{3a})z+(\dfrac{2b^3}{9a^3}-\dfrac{bc}{3a^2}+\dfrac{d}{a})=0"

"H=-\dfrac{b^2}{3a^2}+\dfrac{c}{3a}"

"G=\dfrac{2b^3}{9a^3}-\dfrac{bc}{3a^2}+\dfrac{d}{a}"

"z^3+3Hz+G=0"